OpenStudy (anonymous):

Two pieces of molding 2 and 1/6 inches and 4 and 5/6 inches long are placed end to end. What is the total length of the two pieces of molding together? The total length of the two pieces is inches. Enter your answer as a reduced mixed number or as a whole number.

OpenStudy (anonymous):

2 1/6 + 4 5/6 = 13/6 + 29/6 = 42/6 = 7 The total length of the two pieces is 7 inches.

OpenStudy (mathstudent55):

Above you have the method of converting to fractions. Below you have the method of adding mixed numerals as mixed numerals.